OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). (sin x)(cos x) = 0

OpenStudy (anonymous):

\[\sin x \cos x=0,2\sin x \cos x=0,\sin 2x=0=\sin 0,\sin \pi ,\sin 2 \pi, \sin 3 \pi \] \[2 x=0,\pi,2\pi,3\pi\] find x

OpenStudy (anonymous):

so 0, pi/2, pi, 3pi/2?

OpenStudy (anonymous):

correct.